# On-Chip Frequency Noise Cancellation in Nanomechanical Resonators using Cavity Optomechanics

### v1

@@ Appendix E Optical Non-linear Transduction for difference Frequency generation
We can define the normalized detuning, the cavity coupling efficiency and the normalised amplitude as
$\delta(t)$ $:=\frac{2\Delta(t)}{\kappa}$ (12)
$\bar{\delta}$ $:=\frac{2\bar{\Delta}}{\kappa}$ (13)
$\eta$ $:=\kappa_{\mathrm{ext}}/\kappa$ (14)
$\xi(t)$ $:=\frac{2G_{1}x_{1}}{\kappa}+\frac{2G_{2}x_{2}}{\kappa}$ (15)
Using the input-output relation
@@
$\begin{split}I(t)\approx 1&-\frac{4\eta(1-\eta)}{1+\bar{\delta}^{2}}+\frac{4\eta(1-\eta)2\bar{\delta}}{(1+\bar{\delta}^{2})^{2}}\xi(t)\\ &+\frac{4\eta(1-\eta)(1-3\bar{\delta}^{2})}{(1+\bar{\delta}^{2})^{3}}\xi(t)^{2}\end{split}$ (19)
We define the transduction coefficients to different orders of $\xi$
$D^{(0)}$ $=1-\frac{4\eta(1-\eta)}{1+\bar{\delta}^{2}}$ (20)
$D^{(1)}$ $=\frac{4\eta(1-\eta)2\bar{\delta}}{(1+\bar{\delta}^{2})^{2}}$ (21)
$D^{(2)}$ $=\frac{4\eta(1-\eta)(1-3\bar{\delta}^{2})}{(1+\bar{\delta}^{2})^{3}}$ (22)
The photocurrent is given by
$\begin{split}I(t)=&D^{(0)}+\frac{D^{(2)}}{2}\left(\delta_{1}^{2}+\delta_{2}^{2}\right)+\\ &\frac{D^{(2)}}{2}\delta_{1}\delta_{2}\cos(\Omega_{2}-\Omega_{1})t+\\ &D^{(1)}\delta_{1}\cos\Omega_{1}t+D^{(1)}\delta_{2}\cos\Omega_{2}t+\\ &\frac{D^{(2)}}{2}\left\{\delta_{1}^{2}\cos 2\Omega_{1}t+\delta_{2}^{2}\cos 2\Omega_{2}t+\delta_{1}\delta_{2}\cos(\Omega_{2}+\Omega_{1})t\right\}\end{split}$ (23)
@@
$\begin{split}P_{\mathrm{diff}}&=\frac{{D^{(2)}}^{2}\delta_{1}^{2}\delta_{2}^{2}}{8}\\ &=\frac{32\eta^{2}(1-\eta)^{2}G_{1}^{2}G_{2}^{2}X_{1}^{2}X_{2}^{2}}{\kappa^{2}}\left[\frac{1-3\bar{\delta}^{2}}{(1+\bar{\delta}^{2})^{3}}\right]^{2},\end{split}$ (25)
@@
Figure 9: (Above) Left-half (red-detuned side) of the optical cavity resonance shown in Fig. 4. (Below) Transduced signal for mode $\omega_{1}$ (High Q mode) and the signal at $\omega_{d}$ (Diff signal). the latter shows a decay in signal at the magic detuning, $\bar{\delta}=1/\sqrt{3}$ [38]
@@
Appendix F Data Acquisition and Processing
@@
Appendix H Mass sensing demonstration: Non-linear intrinsic cancellation vs post-processing subtraction
We analyse the data shown in the main-text section IV. Mass Sensing Simulated Experiment using SNR as our figure of merit. Here, we present an alternative quantitative analysis. We refer to the correlations between any two signals $s_{1}(t)$ and $s_{2}(t)$ – we calculate their normalized correlation coefficient $\rho(s_{1},s_{2})=\langle s_{1},s_{2}\rangle/\sqrt{\langle s_{1}\rangle\langle s_{2}\rangle}$, where $\langle$$\rangle$ indicates an average over time, implying that 1 and 0 indicate fully correlated and uncorrelated signals, respectively. We refer to $U(t)$ as the feedback gain that causes an expected frequency shift $\Delta\widetilde{\omega}$. We use the non-averaged data for the below for immediately next discussion where the averaged data used for plotting the main-text Fig. 3 shows very similar numbers.
@@ Appendix H Mass sensing demonstration: Non-linear intrinsic cancellation vs post-processing subtraction
Figure 13: Normalised covariance between the transduced and post-processed $\omega_{d}$ as a function of averaging which translates into an effective sampling rate. The left-most datapoints corresponds to $837\text{\,}\mathrm{S}\mathrm{a}\mathrm{m}\mathrm{p}\mathrm{l}\mathrm{e}\mathrm{s}\mathrm{/}\mathrm{s}$, i.e. without any averaging.

### v2

@@ Appendix F Optical Non-linear Transduction for difference Frequency generation
We can define the normalized detuning, the cavity coupling efficiency and the normalised amplitude as {align} δ(t) := 2Δ(t)κ
¯δ:= 2¯Δκ
η:= κ_ext/κ
ξ
(t) := 2G1x1κ + 2G2x2κ Using the input-output relation
@@
$split I(t)\approx 1&-\frac{4\eta(1-\eta)}{1+\bar{\delta}^{2}}+\frac{4\eta(1-\eta)2\bar{\delta}}{(1+\bar{\delta}^{2})^{2}}\xi(t)\\ &+\frac{4\eta(1-\eta)(1-3\bar{\delta}^{2})}{(1+\bar{\delta}^{2})^{3}}\xi(t)^{2}$ (16)
We define the transduction coefficients to different orders of $\xi$ {align} D^(0) = 1 - 4η(1-η)1 + ¯δ2
D^(1) = 4η(1-η)2¯δ(1+¯δ2)2
D^(2) = 4η(1-η)(1-3¯δ2)(1+¯δ2)3 The photocurrent is given by
$split I(t)=&D^{(0)}+\frac{D^{(2)}}{2}\left(\delta_{1}^{2}+\delta_{2}^{2}\right)+\\ &\frac{D^{(2)}}{2}\delta_{1}\delta_{2}\cos(\Omega_{2}-\Omega_{1})t+\\ &D^{(1)}\delta_{1}\cos\Omega_{1}t+D^{(1)}\delta_{2}\cos\Omega_{2}t+\\ &\frac{D^{(2)}}{2}\left\{\delta_{1}^{2}\cos 2\Omega_{1}t+\delta_{2}^{2}\cos 2\Omega_{2}t+\delta_{1}\delta_{2}\cos(\Omega_{2}+\Omega_{1})t\right\}$ (17)
@@
$split P_{\mathrm{diff}}&=\frac{{D^{(2)}}^{2}\delta_{1}^{2}\delta_{2}^{2}}{8}\\ &=\frac{32\eta^{2}(1-\eta)^{2}G_{1}^{2}G_{2}^{2}X_{1}^{2}X_{2}^{2}}{\kappa^{2}}\left[\frac{1-3\bar{\delta}^{2}}{(1+\bar{\delta}^{2})^{3}}\right]^{2},$ (19)
@@
Figure 11: (Above) Left-half (red-detuned side) of the optical cavity resonance shown in Fig. 3. (Below) Transduced signal for mode $\omega_{1}$ (High Q mode) and the signal at $\omega_{\text{d}}$ (Diff signal). the latter shows a decay in signal at the magic detuning, $\bar{\delta}=1/\sqrt{3}$ [43]
@@
Appendix G Detector Bandwidth
In this section, we characterise the transfer function of our photodetection apparatus. We use a test characterisation setup with an EOM (Exail MPZ-LN-01 Intensity Modulator, operational range up to $3\text{\,}\mathrm{GHz}$) and a polarisation controller to modulate the laser light directly that is incident on the photodiode - transimpedance amplifier detection electronics. The transimpedance amplifier is expected to have an 8 dB smaller signal at $2.16\text{\,}\mathrm{MHz}$ compared to that at 22 $2\text{\,}\mathrm{kHz}$, as per the provided specifications sheet. In order to experimentally validate this, we sweep a modulation tone across a range of $10\text{\,}\mathrm{kHz}$ to $5\text{\,}\mathrm{MHz}$ and record the demodulated signal. We also send a modulation tone at frequencies of interest - $\omega_{\text{d}}\sim$ 22 $2\text{\,}\mathrm{kHz}$, $\omega_{\text{1}}\sim$ $1.06\text{\,}\mathrm{MHz}$ and $\omega_{\text{s}}\sim$ $2.16\text{\,}\mathrm{MHz}$ in successive experiments to plot the voltage Power Spectral Density. We, therefore, experimentally observe over 13 dB of signal drop compared to  $\text{\,}\mathrm{kHz}$ frequency range where the detector response is rather flat. This largely explains our experimental observation in main-text Fig 2(a, d) where the the signal at $\omega_{\text{d}}$ and $\omega_{\text{s}}$ has unequal signal strengths.
Figure 13: Characterization of the responsivity of the detection branch (photodiode and transimpedance amplifier) using an intensity modulator to (a) control the demodulated signal while sweeping the modulation frequency.(b) record the voltage PSD of the laser signal (in grey) and modulate sucessively close to three frequencies of interest, - $\omega_{\text{d}}$, $\omega_{\text{1}}$ and $\omega_{\text{s}}$.
Appendix H Data Acquisition and Processing
@@
Appendix J Sensing demonstration: Non-linear intrinsic cancellation vs post-processing subtraction
We analyse the data shown in the main-text section IV. Sensing Experiment using SNR as our figure of merit. Here, we present an alternative quantitative analysis. We refer to the correlations between any two signals $s_{1}(t)$ and $s_{2}(t)$ – we calculate their normalized correlation coefficient $\rho(s_{1},s_{2})=\langle s_{1},s_{2}\rangle/\sqrt{\langle s_{1}\rangle\langle s_{2}\rangle}$, where $\langle$$\rangle$ indicates an average over time, implying that 1 and 0 indicate fully correlated and uncorrelated signals, respectively. We refer to $U(t)$ as the feedback gain that causes an expected frequency shift $\Delta\widetilde{\omega}$. We use the non-averaged data for the below for immediately next discussion where the averaged data used for plotting the main-text Fig. 3 shows very similar numbers.
@@
Figure 16: Normalised covariance between the transduced and post-processed $\omega_{\text{d}}$ as a function of averaging which translates into an effective sampling rate. The left-most datapoints corresponds to $837\text{\,}\mathrm{S}\mathrm{a}\mathrm{m}\mathrm{p}\mathrm{l}\mathrm{e}\mathrm{s}\mathrm{/}\mathrm{s}$, i.e. without any averaging.
@@
Appendix K Reproducibility on several devices
Several measurements shown in the main text have been repeated over multiple devices to verify the reproducibility of this claim. Below, you find measurements on two such devices. The agreement among all three datasets confirms the reproducibility of the stabilization mechanism and demonstrates that the observed effects are robust across nominally similar devices.
K.1 Device 53
We repeated the full set of measurements on a second optomechanical resonator of identical design. The cavity measurement and detuning Fig. 17(a), thermal PSD Fig. 17(b), effect of optical power on the frequency-PSD Fig. 17(c) are shown. On driving the two modes $\omega_{1}$ and $\omega_{2}$, the voltage PSDs Fig. 18(a-d), the thermal drifts Fig. 18(e-g) and the frequency PSD Fig. 18(h) and Allan deviation analysis Fig. 18(i) are showcased. We also perform the experimental PLL-based frequency-tracking with feedback frequency jumps Fig. 19 to display the same qualitative response as in the primary device, including the characteristic frequency shift.
Figure 17: (a) Normalized reflection trace of the optical mode used in the experiment. Yellow circles: measured reflection as a function of laser wavelength. Yellow dashed line: fit of the optical reflection. Vertical dashed line: cavity detuning used in experiments. (b) Displacement PSD of the calibrated optical frequency noise. The peaks correspond to the thermomechanical noise stemming from the perimeter modes of the polygon resonator. Modes marked with red and blue are used in the experiment. (c) Mechanical frequency noise PSD of the high-$Q$ mode at different optical powers.
Figure 18: Properties of signal at $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$. Displacement PSD of demodulated signal around (a) $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$ and (c) $\omega_{s}=\omega_{1}+\omega_{2}$. The mechanical modes at $\omega_{1}$ (red) and $\omega_{2}$ (blue) are driven with separate PLLs. (b) Thermal mechanical spectrum without external drive, same as in Fig. 17 (b) to highlight the modes $\omega_{1}$ and $\omega_{2}$. (d)-(f) Spectograms of the difference signal and the displacements of the OOP and the IP modes. The color bar corresponds to the normalized PSD. (g) Double-sided frequency-PSD for $\omega_{1}$, $\omega_{2}$, and $\omega_{\mathrm{d}}$ spanning $150\text{\,}\mathrm{s}$ with a data rate of $14.3\text{\,}\mathrm{kHz}$. We use the PLLs for estimating $\omega_{1}$ and $\omega_{2}$, while $\omega_{\mathrm{d}}$ is calculated using phase-to-frequency conversion [42]. (h) Absolute Allan deviation $\sigma$ in  $\text{\,}\mathrm{Hz}$ calculated using the same traces as in (h).
K.2 Device 26 on a different chip
We characterise yet another device on a different chip with an identical design but different dimensions. The characterisation of the device is provided in Fig. 20. The improved performance at $\omega_{\text{d}}$ is showcased in Fig. 21.
Figure 20: (a) Normalized reflection trace of the optical mode used in the experiment. Yellow circles: measured reflection as a function of laser wavelength. Yellow dashed line: fit of the optical reflection. Vertical dashed line: cavity detuning used in experiments. (b) Displacement PSD of the calibrated optical frequency noise. The peaks correspond to the thermomechanical noise stemming from the perimeter modes of the polygon resonator. Modes marked with red and blue are used in the experiment. (c) Mechanical frequency noise PSD of the high-$Q$ mode at different optical powers.
Figure 21: Properties of signal at $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$. Displacement PSD of demodulated signal around (a) $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$ and (c) $\omega_{s}=\omega_{1}+\omega_{2}$. The mechanical modes at $\omega_{1}$ (red) and $\omega_{2}$ (blue) are driven with separate PLLs. (b) Thermal mechanical spectrum without external drive, same as in Fig. 20 (b) to highlight the modes $\omega_{1}$ and $\omega_{2}$. (d)-(f) Spectograms of the difference signal and the displacements of the OOP and the IP modes. The color bar corresponds to the normalized PSD. (g) Double-sided frequency-PSD for $\omega_{1}$, $\omega_{2}$, and $\omega_{\mathrm{d}}$ spanning $150\text{\,}\mathrm{s}$ with a data rate of $14.3\text{\,}\mathrm{kHz}$. We use the PLLs for estimating $\omega_{1}$ and $\omega_{2}$, while $\omega_{\mathrm{d}}$ is calculated using phase-to-frequency conversion [42]. (h) Absolute Allan deviation $\sigma$ in  $\text{\,}\mathrm{Hz}$ calculated using the same traces as in (h).